# Languages Transferred Within the Encoder: On Representation Transfer in Zero-Shot Multilingual Translation

### v1

@@ Abstract
Understanding representation transfer in multilingual neural machine translation can reveal the representational issue causing the zero-shot translation deficiency. In this work, we introduce the identity pair, a sentence translated into itself, to address the lack of the base measure in multilingual investigations, as the identity pair represents the optimal state of representation among any language transfers. In our analysis, we demonstrate that the encoder transfers the source language to the representational subspace of the target language instead of the language-agnostic state. Thus, the zero-shot translation deficiency arises because representations are entangled with other languages and are not transferred effectively to the target language. Based on our findings, we propose two methods: 1) low-rank language-specific embedding at the encoder, and 2) language-specific contrastive learning of the representation at the decoder. The experimental results on Europarl-15, TED-19, and OPUS-100 datasets show that our methods substantially enhance the performance of zero-shot translations by improving language transfer capacity, thereby providing practical evidence to support our conclusions.111Partial works done during Zhi Qu’s internship at ASTREC of NICT, Japan. Our codes are available at https://github.com/zhiqu22/ZeroTrans.
@@
Zhi Qu† Chenchen Ding†‡ Taro Watanabe† †Nara Institute of Science and Technology, Japan {qu.zhi.pv5, taro}@is.naist.jp ‡National Institute of Information and Communications Technology, Japan chenchen.ding@nict.go.jp
@@ 1 Introduction
State-of-the-art neural machine translation systems are adaptable to multilingualism, resulting in a single encoder-decoder model that executes arbitrary translations by adding a tag specified to the target language at the beginning of source sentence Firat et al. (2016); Johnson et al. (2017); Wu et al. (2021). Multilingual neural machine translation (MNMT) is theoretically attractive because zero-shot translations, i.e., translations unseen in training, allow the training of a multilingual model with minimal cost. Unfortunately, the performance of zero-shot translations always lags behind Aharoni et al. (2019); Arivazhagan et al. (2019a); Gu et al. (2019); Yang et al. (2021); Pan et al. (2021); Chen et al. (2023).
Representational analysis in MNMT models guides the improvement of zero-shot translation. However, two different perspectives are demonstrated by prior works: 1) the encoder clusters the different translation representations by the target language Kudugunta et al. (2019); Liu et al. (2021), as shown in Figure 1(a); 2) an ideal encoder is considered to learn only language-agnostic representations, where general cross-lingual features are modeled and transferred among languages Pan et al. (2021); Gu and Feng (2022); Gao et al. (2023), as shown in Figure 1(b). In this work, our investigation proves that these two perspectives are two sides of the same coin, as the encoder aligns translation representations from different source languages within the subspace of the target language.
We introduce the identity pair, which is a pseudo pair translating a sentence to itself, as a proxy for the optimal representation of a language. We then use multiple analytical methods, including singular value canonical correlation analysis (SVCCA) Raghu et al. (2017) and t-distributed stochastic neighbor embedding (t-SNE) van der Maaten and Hinton (2008), to systematically analyze and quantify the representation transfer in MNMT models. We show that transferring representations from the source language to the target language is a major task of the encoder and is entirely accomplished by the encoder. Therefore, the zero-shot translation deficiency arises from the translation representations being entangled with other languages rather than being effectively transferred to the expected target language within the encoder.
@@ 2.1 Multilingual Neural Machine Translation
2.2 Prior Investigations of MNMT
Two tools are used in prior investigations: 1) The t-distributed stochastic neighbor embedding (t-SNE) van der Maaten and Hinton (2008) converts similarities between a set of vector representations to joint probabilities; 2) The singular value canonical correlation analysis (SVCCA) Raghu et al. (2017) compares the similarity between two vector representations222We follow Liu et al. (2021) to measure SVCCA scores at the sentence level. Appendix A shows how to compute SVCCA scores.. The visualization using t-SNE shows that translation representations, which belong to different source languages but with the same semantics, are aligned at the output of the encoder Pan et al. (2021); Gao et al. (2023), suggesting the ideal encoder is language-agnostic. Specifically, translation representations from different source languages are transferred to a language-agnostic state before being transferred to the target language. On the contrary, the investigation of Kudugunta et al. (2019) based on SVCCA demonstrates that translation representations possess the features of the target language. However, the results derived from measuring SVCCA scores between real translation pairs are disturbed by the lack of base measures. For instance, the translation from English to German, denoted by ende, cannot serve as a base measure for the translation from another language xde, because the representation of ende is not equivalent to the representation of de.
@@ 3 Investigating Representation Transfer in MNMT
We conduct preliminary experiments to investigate representations by introducing identity pairs as base measures. Two different datasets, Europarl-15 Koehn et al. (2005); Dabre and Kurohashi (2019) and TED-19 Ye et al. (2018), are applied to verify whether our observations are general. Appendix B introduces the detailed information of those two datasets. Then, following Kudugunta et al. (2019), our investigation is based on Transformer models with 6 encoder and decoder layers, where Appendix C introduces the detailed model settings.
@@ 3.1 Identity Pairs
Identity pairs refer to the translation pairs translating from themselves into themselves to present the optimal state of processing language features, i.e., the semantics and syntax of the source sentence by the model. On the encoder side, we derive the representation from a language translating to itself, i.e., , where is the language tag of , with the aim of recovering the source sentence from the hidden representations without inference on the decoder side. We derive the representation in the decoder from the gold translation of .
@@ 3.2 Language Transfer Within the Encoder
Given two languages ① and ②, two sentences of ① and of ② are grouped by three cases when comparing the representations of the encoder: (i) comparing and to show how target language features are encoded333 indicates the identity of ①, i.e., ①①, and indicates a sentence of ② translating to the sentence of ① instructed by the language tag of ①, i.e., ②①.; (ii) comparing and to show how source language features are encoded; (iii) comparing two different identities, and .
The two models trained by Europarl-15 and TED-19 show the same tendency in Figure 2, that is, the language features for ① of (i) consistently increase in both cases involving the central language in Figures 2(a) and 2(c), and non-central languages in Figures 2(b) and 2(d). The target language feature of ① emerges as the primary factor that affects representations at the fifth and sixth layers when the cases of (i), (ii), and (iii) are compared. Therefore, we can conclude that the language features of the representations are transferred to the target side within the encoder. Meanwhile, we also observe that the scores of (iii) are close to or even exceed those of (ii) at some layers both in Figure 2. This proves that the feature of the source language is not the primary factor for representations in the encoder because representations are transferred to the subspaces of target languages. Thus, the comparison between (ii) and (iii) supports the conclusion that language transfer is completed within the encoder.
On the other side, identity pairs also allow the measurement of the alignment of different languages in the target language space through t-SNE. Compared with the sentence-level measurement of Pan et al. (2021); Gao et al. (2023), we measure the alignment of representations at the token level. As shown in Figure 3, semantic similarity causes the representations to cluster together. Moreover, as shown in Appendix H, these representations are not clustered before being processed by the encoder, and the case with different target languages has a higher overall variance. Combined with the finding that the encoder transfers the representation of the source language to the target language, the evidence further suggests that there is no general and cross-lingual state for directly sharing semantic information within the encoder, and the alignment shown in Figure 1(b) occurs in the representational subspace of the target language.
@@ 3.3 Entanglements Hindering the Transfer
Although the investigation in Section 3.2 shows that the representations gradually transfer to the target language in a translation pair, the representation spaces of multiple languages may potentially entangle with each other, resulting in the failure of the zero-shot translation Qu and Watanabe (2022). To further illustrate the relationship between different languages, we use t-SNE and BiKDE to visualize the representations at the output of the encoder for the several identity pairs in Europarl-15 and TED-19. Figures 4(a) and 4(d) show that different identity pairs are uniformly distributed in the representational space. This distribution proves again that the encoder is language-specific because each language has an isolated representational subspace.
Compared with identity pairs that represent the ideal capability of the model in processing languages, the distributions plotted in Figures 4(b) and 4(e) reflect the actual capacity for the supervised translation of enx. Figures 4(b) and 4(e) show the distribution of representations in the pairs translating from en, which are similar to that of identity pairs. The difference between identity pairs and supervised language pairs can be attributed to the influence of the source language information, which hinders the full use of the target language information learned by the encoder.
Moreover, the language-specific subspaces cannot be clearly separated for zero-shot translations, as shown in Figures 4(c) and 4(f). Specifically, all representations are entangled around the supervised language pair of xen, which hinders these representations from transferring into the ideal subspaces of the target language. This aligns with Qu and Watanabe (2022) that non-central languages are entangled with the central language, i.e., English, which explains the weakness of zero-shot translation compared with supervised translation, namely, the focal point for improving the transferability of representations is to improve the extent of language transfer in the encoder.
@@ 3.4 Language Features in the Decoder
The decoder works on representations with target language features because the source language has already been transferred into the target language subspace (Section 3.2). Therefore, representational entanglements (Section 3.3) also affect the decoder. We further investigate the importance of target language features versus semantics in the decoder. Given two sentences of language ③ and of language ④, the decoder representation of is considered as the base measure. We group two cases: (iv) For each sentence in a test set, we identify the pair with the lowest SVCCA score in the encoder representation to derive a that is distant from . Then, we compare it with the base measure to show the importance of target language features; (v) We compare the base measure and to show the importance of semantics. The two scenarios shown in Figure 5 present the same trend, which is that (iv) maintains high scores despite their semantics being entirely different. At the top layers of the decoder, the gradually increasing difference between (iv) and (v) confirms that the decoder tends to learn the target language specificity Sen et al. (2019). However, Figure 5 shows that, for (iv), a wider interquartile range exists at the bottom layers of the decoder, and its scores are close to those of (v), which implies the weakness of the distinguishing languages in zero-shot translations.
@@ 5.1 Setup
Our experiments comprise three English-centric scenarios, i.e., the training and validation sets only involving translation pairs translating to en or from en. 1) Europarl-15 is a sub-collection of MMCR4NLP Koehn et al. (2005); Dabre and Kurohashi (2019), resulting in 5.3 million training translation pairs. 2) TED-19 collects 19 languages from TED Talks Ye et al. (2018), where the training set comprises 6.5 million pairs in total. 3) OPUS-100 Zhang et al. (2020a); Yang et al. (2021) consists of 95 languages and 109.2 million training pairs. Generally, the validation and test sets of a pair contain 2,000 pairs. More detailed information of those datasets can be found in Appendix B.
| Europarl-15 | TED-19 | OPUS-100 | ||||||||||||||
| BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | |||||||||||
| SemAli | 37.27 | 43.06 | 25.25 | 88.42 | 95.69 | 84.43 | 23.55 | 28.67 | 13.45 | 83.43 | 93.36 | 78.91 | 22.35 | 28.29 | 6.42 | 72.00 |
| LoLE | 37.62 | 43.50 | 26.09 | 88.51 | 95.72 | 84.81 | 24.39 | 29.72 | 13.20 | 83.74 | 93.54 | 78.65 | 23.15 | 28.28 | 7.92† | 73.32† |
| BOTH | 37.67† | 43.51 | 26.20† | 88.50 | 95.72 | 84.85† | 24.49 | 29.79 | 13.31 | 83.76 | 93.56 | 78.76 | 23.40 | 28.27 | 7.97† | 73.10† |
@@
In experiments of training from scratch, we implement a Transformer model with 6 encoder and decoder layers. Then, three pre-trained models444Notably, those models are trained by adding a source language tag at the encoder and a target language tag at the decoder. In fine-tuning, we keep their strategy, because it is not conflict with adding extra target information by LoLE. are utilized in our experiments of fine-tuning, including M2M-418M, M2M-1.2B Fan et al. (2020) and mBART50 Liu et al. (2020). Based on the ablation study conducted in the validation set in Europarl-15 and TED-19 shown in Appendix F, we apply LoLE in the fifth encoder layer, set to 128, and set and of LCLR to 64 and 30, respectively. We enlarge to 256 and to 128 for models trained on OPUS-100 and three pre-trained models because they involve more languages. When we solely apply LCLR, we set the position to the bottom decoder layer based on the findings in Section 3.4. When we integrate both LoLE and LCLR into a model, we relocate LCLR to the second-bottom decoder layer because of the improved language features of the encoder representations. More detailed settings of models and training details can be found in Appendix C.
Baselines and Evaluation
The vanilla Transformer Vaswani et al. (2017); Johnson et al. (2017) is a baseline in experiments, denoted by Vanilla. Additionally, three representative methods are reproduced in experiments of training from scratch with our datasets and settings: 1) SemAli Pan et al. (2021), which is a state-of-the-art MNMT model without fine-tuning that uses contrastive learning in the encoder output for semantic alignment; 2) DisPos Liu et al. (2021), which removes one residual connection in the encoder to weaken the source language information; and 3) TLP Yang et al. (2021), which has an extra loss introduced to identify the language of the decoder’s output representation. In the evaluation, we set the beam size to 4 in inference. We not only use SacreBLEU Papineni et al. (2002); Post (2018) to evaluate the quality of inferences but also report BERTScore Zhang et al. (2020b) of inferences at the representation level. We also conduct statistical significance testing Koehn (2004) using paired bootstrap resampling with 1,000 iterations and 0.5 resampling ratio, consequently, the case of 0.05 means that the difference is significant.

### v2

@@ Abstract
Understanding representation transfer in multilingual neural machine translation (MNMT) can reveal the reason for the zero-shot translation deficiency. In this work, we systematically analyze the representational issue of MNMT models. We first introduce the identity pair, translating a sentence to itself, to address the lack of the base measure in multilingual investigations, as the identity pair can reflect the representation of a language within the model. Then, we demonstrate that the encoder transfers the source language to the representational subspace of the target language instead of the language-agnostic state. Thus, the zero-shot translation deficiency arises because the representation of a translation is entangled with other languages and not transferred to the target language effectively. Based on our findings, we propose two methods: 1) low-rank language-specific embedding at the encoder, and 2) language-specific contrastive learning of the representation at the decoder. The experimental results on Europarl-15, TED-19, and OPUS-100 datasets show that our methods substantially enhance the performance of zero-shot translations without sacrifices in supervised directions by improving language transfer capacity, thereby providing practical evidence to support our conclusions. Codes are available at https://github.com/zhiqu22/ZeroTrans.
@@
Zhi Qu††thanks: This work was done during the first author’s internship at Advanced Speech Translation Research and Development Promotion Center, National Institute of Information and Communications Technology, Kyoto, Japan.† Chenchen Ding†‡ Taro Watanabe† †Nara Institute of Science and Technology, Japan {qu.zhi.pv5, taro}@is.naist.jp ‡National Institute of Information and Communications Technology, Japan chenchen.ding@nict.go.jp
@@ 1 Introduction
State-of-the-art neural machine translation systems are adaptable to multilingualism, resulting in a single encoder-decoder model that executes arbitrary translations by adding a tag specified to the target language at the beginning of source sentence Firat et al. (2016); Johnson et al. (2017); Wu et al. (2021). Multilingual neural machine translation (MNMT) is theoretically attractive because zero-shot translations, i.e., translations unseen in training, allow the training of a multilingual model with minimal cost. Unfortunately, the performance of zero-shot translations always lags behind Aharoni et al. (2019); Arivazhagan et al. (2019a); Gu et al. (2019); Yang et al. (2021); Pan et al. (2021); Chen et al. (2023a).
Representational analysis in MNMT models can guide the improvement of zero-shot translation. However, two contrary opinions are demonstrated by prior works: (1) the encoder clusters translation representations based on the target language Kudugunta et al. (2019); Liu et al. (2021); Tan and Monz (2023); Stap et al. (2023); Sun et al. (2024), as illustrated in Figure 1(a); (2) an ideal encoder is expected to learn language-agnostic representations, capturing general cross-lingual features that are transferable across languages Pan et al. (2021); Gu and Feng (2022); Gao et al. (2023); Bu et al. (2024), as shown in Figure 1(b). In this work, we aim to analyze and reconcile this discrepancy. We first introduce the identity pair, a pseudo pair translating a sentence to itself. Specifically, the analyses conducted by those prior works rely on real translation pairs, leading to inaccurate results, as a translation pair cannot serve as a base measure for another pair. The identity pair, however, addresses this issue by serving as a proxy for the optimal representation of a language instead of a translation pair. Then, multiple analytical methods are employed to show the representation transfer within MNMT models. Our findings offer a unified perspective on these two opinions: the encoder transfers translation representations into the target language subspace, where different source languages are semantically aligned. Thus, the zero-shot translation deficiency stems from the failure to transfer the translation representation to the target language, as it becomes entangled with representations of other languages in the encoder.
@@ 2.1 Multilingual Neural Machine Translation
2.2 The Discrepancy in Prior Works
Pan et al. (2021); Gao et al. (2023); Bu et al. (2024) state that, for an encoder-decoder MNMT model, an ideal encoder is regarded as transferring the source sentence into a language-agnostic state, preserving only semantic information.111Although Pan et al. (2021) proposed that the ideal output of the encoder is language-agnostic by adding a source language tag at the beginning of the encoder, the follow-up works Gao et al. (2023); Bu et al. (2024) practiced this concept with adding a target language tag, which is the main strategy investigated in this work. As evidence, the t-distributed stochastic neighbor embedding (t-SNE) van der Maaten and Hinton (2008), which can convert similarities between vectors into joint probabilities, has been used to show that representations of sentences from different languages are aligned at the output of the encoder when sharing the same semantics. However, this result contrasts with the findings of Kudugunta et al. (2019); Stap et al. (2023); Tan and Monz (2023). Specifically, using the singular value canonical correlation analysis (SVCCA) Raghu et al. (2017) to compare the similarity between two vectors, i.e., the sentence representations of two translations, reveals that the encoder tends to transfer the representation into a state with target language features.
We argue that this discrepancy stems from the lack of a base measure. Namely, those works always compare the representations of real translation pairs in which different analysis methods lead to different results. For instance, the translation from English to German, denoted by ende, cannot be accurately measured by comparing it with another translation from a different language xde, because ende is expected to be measured by the language representation of either de or en. Thus, proposing a base measure is necessary to draw the same conclusion from different analysis methods, e.g., t-SNE or SVCCA222We follow Liu et al. (2021) to measure SVCCA scores at the sentence level, which is introduced in Appendix A..
@@ 3 Investigating Representation Transfer in MNMT
We conduct preliminary experiments to investigate representations by introducing identity pairs as base measures using two different datasets, Europarl-15 Koehn et al. (2005) and TED-19 Ye et al. (2018), which are introduced in Appendix B in detail. Then, following Kudugunta et al. (2019), our investigation is based on Transformer models with 6 encoder and decoder layers. We also investigate scenarios with 8 and 10 encoder layers. Appendix C introduces the detailed model settings.
@@ 3.1 Identity Pairs
An identity pair refers to a sentence pair translating from one sentence into itself to represent the optimal state of processing language features, i.e., the semantics and syntax of the source sentence by the model. Notably, our models are only trained by translating from one language to another. In this setup, the identity pair is a zero-shot translation, which does not simply copy the input to the output.333This claim is supported by Qu and Watanabe (2022), which demonstrate another zero-shot scenario: removing the language tag during inference results in any source sentence being translated into English. Thus, the identity pair indeed presents a translation process by adding a language tag. On the encoder side, we derive the representation from a language translating to itself, i.e., , where is the language tag of , with the aim of recovering the source sentence from the hidden representations without inference on the decoder side. We also derive the representation in the decoder from the gold translation of .
@@ 3.2 Language Transfer Within the Encoder
Given two languages ① and ②, we follow Pan et al. (2021); Liu et al. (2021) to obtain sentence-level representations for in ① and in ② by applying mean pooling over token representations. We then organize the comparisons into three cases to analyze the variation in encoder representations: (i) comparing and to show how target language features are encoded; (ii) comparing and to show how source language features are encoded; (iii) comparing two different identities, and 444 indicates the identity of ①, i.e., ①①, and indicates a sentence of ② translating to the sentence of ① instructed by the language tag of ①, i.e., ②①..
The two models trained by Europarl-15 and TED-19 show the same tendency in Figure 2, i.e., the language features for ① of (i) consistently increase in both cases involving the central language, i.e., English, in Figures 2(a) and 2(c), and non-central languages in Figures 2(b) and 2(d). The target language feature of ① emerges as the primary factor that affects representations at the fifth and sixth layers when the cases of (i), (ii), and (iii) are compared. Therefore, we can conclude that the language features of the representations are transferred to the target side within the encoder. Meanwhile, we observe that the scores of (iii) are close to or even exceed those of (ii) at some layers both in Figure 2. This proves that the feature of the source language is not the primary factor for encoding representations because representations are transferred to the subspaces of target languages. Thus, the comparison between (ii) and (iii) supports that language transfer is completed within the encoder.
To validate the generality of this conclusion, we extend our analysis to models with 8 and 10 encoder layers (Figure 3). The same trends hold: (i) continues to show increasing similarity scores, with the final values even higher than in the 6-layer setting, suggesting stronger target language alignment. Again, the relationship between (ii) and (iii) remains consistent, further confirming that the source language is not the dominant factor in shaping encoder representations. These results demonstrate that language transfer within the encoder is robust across different architectural depths.
On the other hand, identity pairs also allow the measurement of the alignment of different languages in the target language space through t-SNE. Compared with the sentence-level measurement of Pan et al. (2021); Gao et al. (2023), we measure the alignment of representations at the token level. As shown in Figure 4, semantic similarity causes the representations to cluster together. Moreover, as shown in Appendix G, these representations are not clustered before being processed by the encoder, and the case with different target languages has a higher overall variance. Combined with the finding that the encoder transfers the representation of the source language to the target language, the evidence further suggests that there is no general and cross-lingual state for directly sharing semantic information within the encoder, and the alignment shown in Figure 1(b) occurs in the representational subspace of the target language.
@@ 3.3 Entanglements Hindering the Transfer
Although the investigation in Section 3.2 shows that the representations gradually transfer to the target language in a translation pair, the representation spaces of multiple languages may potentially entangle with each other, resulting in the failure of the zero-shot translation Qu and Watanabe (2022). To further illustrate the relationship between different languages, we use t-SNE and BiKDE to visualize the representations at the output of the encoder for the several identity pairs in Europarl-15 and TED-19. Figures 5(a) and 5(d) show that different identity pairs are uniformly distributed in the representational space. This distribution proves again that the encoder is language-specific because each language has an isolated representational subspace.
Compared with identity pairs that represent the ideal capability of the model in processing languages, the distributions plotted in Figures 5(b) and 5(e) reflect the actual capacity for the supervised translation of enx. Figures 5(b) and 5(e) show the distribution of representations in the pairs translating from en, which are similar to that of identity pairs. The difference between identity pairs and supervised language pairs can be attributed to the influence of the source language information, which hinders the full use of the target language information learned by the encoder.
Moreover, the language-specific subspaces cannot be clearly separated for zero-shot translations, as shown in Figures 5(c) and 5(f). Specifically, all representations are entangled around the supervised language pair of xen, which hinders these representations from transferring into the ideal subspaces of the target language. This aligns with Qu and Watanabe (2022) and Stap et al. (2023) that multilingual representations are entangled, which explains the weakness of zero-shot translation compared with supervised translation, suggesting that improving the transferability of representations is attributed to the extent of language transfer within the encoder.
@@ 3.4 Language Features in the Decoder
We further investigate the importance of target language features versus semantics in the decoder. Given two sentences of language ③ and of language ④, the decoder representation of is considered as the base measure. We group two cases: (iv) For each sentence in a test set, we identify the pair with the lowest SVCCA score in the encoder representation to derive a that is distant from . Then, we compare it with the base measure to show the importance of target language features; (v) We compare the base measure and to show the importance of semantics. The two scenarios shown in Figure 6 present the same trend, which is that (iv) maintains high scores despite their semantics being entirely different. At the top layers of the decoder, the gradually increasing difference between (iv) and (v) confirms that the decoder tends to learn the target language specificity Sen et al. (2019). However, Figure 6 shows that, for (iv), a wider interquartile range exists at the bottom layers of the decoder, and its scores are close to those of (v), which implies the weakness in distinguishing languages for zero-shot translations.
@@ 4.1 Low-Rank Embedding for the Encoder
Europarl-15 | TED-19 | OPUS-100 | ||||||||||||||
BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | |||||||||||
SemAli | 37.27 | 43.06 | 25.25 | 88.42 | 95.69 | 84.43 | 23.55 | 28.67 | 13.45† | 83.43 | 93.36 | 78.91† | 22.35 | 28.29 | 6.42 | 72.00 |
LoLE | 37.62 | 43.50 | 26.09† | 88.51 | 95.72 | 84.81 | 24.39 | 29.72 | 13.20 | 83.74 | 93.54 | 78.65 | 23.15 | 28.28 | 7.92† | 73.32† |
BOTH | 37.67 | 43.51 | 26.20† | 88.50 | 95.72 | 84.85† | 24.49 | 29.79 | 13.31† | 83.76 | 93.56 | 78.76† | 23.40 | 28.27 | 7.97† | 73.10† |
@@ 5.1 Setup
Our experiments comprise three popular English-centric datasets, i.e., the training and validation sets only involving translation pairs translating to en or from en, including Europarl-15 Koehn et al. (2005); Dabre and Kurohashi (2019), TED-19 Ye et al. (2018) and OPUS-100 Zhang et al. (2020a); Yang et al. (2021). The details of those datasets can be found in Appendix B.
Evaluation
We evaluate the performance of models on the test sets of those three datasets and set the beam size to 4 in inference. We employ SacreBLEU Papineni et al. (2002); Post (2018) to evaluate the quality of inferences at the word level and report BERTScore Zhang et al. (2020b) of inferences at the representation level. We measure the off-target ratio on zero-shot translations as a supplement. We also conduct statistical significance testing Koehn (2004). We describe our motivation in selecting evaluation metrics, the evaluation details, and the implementation of statistical significance testing in Appendix H.
@@
When training from scratch, we implement a Transformer model with 6 encoder and decoder layers. Given that those three datasets have different sizes, we set different hyper-parameters in training. Then, three open-source models555Note that, those models are trained by adding a source language tag at the encoder and a target language tag at the decoder. In fine-tuning, we keep the original strategy. are utilized in fine-tuning experiments, including M2M-418M, M2M-1.2B Fan et al. (2020) and mBART50 Tang et al. (2020). The hyper-parameter settings can be found in Appendix C. Additionally, hyper-parameters are selected based on the ablation studies conducted on the validation sets, which is reported in Appendix E.
Baselines
Vanilla Transformer Vaswani et al. (2017); Johnson et al. (2017) is one of the baselines, denoted by Vanilla in the experiments of training from scratch. Then, the baseline in fine-tuning experiments is the full-parameter fine-tuning, denoted by F.T.. Moreover, three representative methods are reproduced in our experiments of training from scratch, the standard of baseline selection is as follows:
SemAli: Pan et al. (2021) think the encoder output is language-agnostic, so they align the semantic information across different languages at the output of the encoder. However, our analysis shows that this viewpoint is inaccurate because the semantic information is aligned by the subspace of the target language instead of the real language-agnostic. When there are not any additional parameters introduced, SemAli still is the de-facto SOTA based on regularizing representations in MNMT.
5.2 Results
First of all, Gu et al. (2019); Liu et al. (2021) pointed out that the vanilla Transformer is superior in supervised translation directions, i.e., en<->x, because the model excessively focuses on English, which is the language dominating the training set, to lose its generalization on non-English languages, i.e., the zero-shot translation. Moreover, Huang et al. (2023); Chen et al. (2023b) showed that improving the zero-shot may come at the expense of supervised performance. In this work, our methods significantly improve the zero-shot translation without degrading the supervised performance in both training from scratch and fine-tuning.
Table 1 shows the experimental results of training from scratch. In supervised translations of Europarl-15/TED-19/OPUS-100, LoLE shows divergent results of 0.13/-0.04/-0.22 on enx and 0.11/0.05/-0.02 on xen. Similarly, LCLR shows diverse results of -0.05/-0.01/-0.03 and 0.04/-0.01/0.07, respectively. Then, Both achieves the results of 0.18/-0.02/0.03 on enx and 0.12/0.12/-0.03 on xen. In zero-shot translations, BOTH outperforms Vanilla 1.55/1.33/2.93 for BLEU and 0.58/1.02/3.12 for BERTScore. Our models perform best in zero-shot translations of Europarl-15 and OPUS-100, and the improvements in zero-shot translations are always statistically significant. Note that, although SemAli achieves the best zero-shot translation performance in TED-19, the supervised performance of SemAli is significantly degraded compared to Vanilla, which is a common and unresolved problem Gu et al. (2019); Zhang et al. (2020a); Liu et al. (2021). On the contrary, our methods not only perform competitively with SemAli in zero-shot translations but also benefit the supervised translation capacity. Moreover, these two proposed methods are orthogonal, which can be proved by assessing LoLE, LCLR and Both individually: (1) LoLE achieves gains of 1.53/1.22/2.85 for BLEU and 0.54/0.91/3.34 for BERTScore; (2) LCLR improves 1.06/0.14/0.09 and 0.39/0.13/0.06 scores; (3) The gains of Both are always higher than LoLE and LCLR. In addition, we can observe that the improvement of LCLR is limited in TED-19 and OPUS-100, which can be attributed to the diverse languages involving in these two datasets and being easily distinguished by the vanilla decoder. This result also supports that the main challenge of MNMT is the transfer within the encoder. Thus, we can conclude that our methods substantially benefit the zero-shot translation capacity of MNMT models.